# Self interacting scalar field theory in general curved spacetimes at zero and finite temperature revisited

### v1

@@ Abstract
We revisit the problem of spontaneous symmetry breaking (SSB), its restoration and phase transition for a self interacting quantum scalar field in a general curved background, at zero and finite temperature. To the best of our knowledge, most of the earlier computations in this context have been done in the linear order in curvature, which may not be very suitable for the Ricci flat spacetimes. One of our objectives is to see whether the higher order terms can bring in qualitatively new physical effects, and thereby attempting to fill in this gap in the literature. We use the Bunch-Parker local momentum space representation of the Feynman propagator. We compute the renormalised, background spacetime curvature (up to quadratic order) and temperature dependent one loop effective potential for $\phi^{4}$ plus $\phi^{3}$ self interaction. In particular for the de Sitter spacetime, we have shown for $\phi^{4}$-theory that we can have SSB for a scalar even with a positive rest mass squared and non-minimal coupling, at zero temperature. This cannot be achieved by the linear curvature term alone and the result remains valid for a very large range of renormalisation scale. For a phase transition, we have computed the leading curvature correction to the critical temperature. At finite temperature, symmetry restoration is also demonstrated. We also extend some of the above results to two loop level. The symmetry breaking in de Sitter at two loop remains present. We have further motivated the necessity of treating this problem non-perturbatively in some instances.
@@ 1 Introduction
The motivation of studying quantum field theory in curved backgrounds is to investigate the effects of spacetime curvature or geometry on the fluctuations or propagation of quantum fields living in it. This much cultivated yet still active subject, originally supposed to be a prelude to a full theory of quantum gravity, has led to interesting physical predictions like the particle creation in cosmological or black hole backgrounds, fluctuations in the early inflationary universe, regularisation of the stress energy momentum tensor and so on, over decades. We refer our reader to [1, 2, 3, 4] and references therein for a vast historic overview.
@@
We are interested to look into a self interacting scalar field theory at zero as well as at finite temperature in a general, regular curved background in this paper. Thermal fields in a curved background may exist in various different scenarios. For instance, the thermal fields living in the early universe, or the hot plasma residing in a star or accreting onto a black hole, the blackbody radiation emitted by a black hole, and so on.
@@
An interesting aspect of quantum field theory with a background or classical scalar field is the existence of an effective potential. Such potential arises due to the loop integration over the quantum fluctuations and it adds up to the tree level self interaction. The total effective potential (i.e., the tree plus the loop correction) can be qualitatively different in shape from that of the tree level. Such effective potential may lead to the highly interesting physical phenomenon of spontaneous symmetry breaking, developed in the seminal works in [5, 6, 7, 8]. We further refer our reader to [9, 10, 11, 12] and references therein for various aspects associated with symmetry breaking and effective potentials. Computation of effective potentials for field theories at finite temperature is also a physically well motivated problem. Interestingly, the effective potential at high temperature ceases to have any symmetry breaking feature even if at zero temperature it has so, known as the symmetry restoration, e.g. [7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. We further refer our reader to [25, 26, 27] and references therein for vast discussion on these issues.
@@
We do not have the scope here to review in detail the developments of quantum field theory in curved spacetimes, and the effective action formalism. We refer our reader to [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] and references therein for some aspects of propagators, renormalisation, effective action and symmetry breaking issues at zero temperature. For thermal propagators and finite temperature effects in curved spacetime, we refer our reader to [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] and also references therein. As far as symmetry breaking and phase transition in a general curved background is concerned, most of the earlier works use linear in the curvature approximation and then employ some renormalisation group technique for resummation (e.g. [33, 35, 38, 40] and also references therein). For example in [38], a renormalisation group derivation of the effective potential for a massless scalar coupled to gauge field, and discussion on phase transition can be seen. Even though this is a powerful method, it does not work for the Ricci flat spacetimes. In particular, the $R\phi^{2}$-type non-minimal coupling does not contribute to the effective potential for such spacetimes. At finite temperature on the other hand, the Page approximation technique can be useful for static spacetimes [62]. To the best of our knowledge, the thermal effective Lagrangian density in a general curved spacetime for $\phi^{4}$ theory, up to quadratic order in the curvature was first derived in [67], by employing a proper time expansion technique.
@@
The rest of the paper is organised as follows. In the next section, we discuss very briefly the basic technical framework we work in. In 3, we briefly derive the one loop effective potential at zero temperature, up to the quadratic order of the spacetime curvature for quartic and cubic self interactions. In 4, we extend these results to finite temperature, with high temperature approximation. Although as we mentioned, the above problem was first addressed in [67], our notation and the choice of the renormalisation scale will be a bit different. Due to this reason, we shall keep briefly the computations, for the sake of completeness and clarity. We discuss spontaneous symmetry breaking, its restoration and phase transition at zero and finite temperatures in 5, 5.1. We next generalise some of these results to two loop in 6. In particular, the contribution from the two loop double bubble diagram under local approximation to the effective potential, at finite temperature has been explicitly presented, 6.1. Finally we conclude in 7 with some future motivation for non-perturbative calculations. For the de Sitter spacetime in particular, we have shown that we can have symmetry breaking for a scalar even with positive rest mass squared and a positive non-minimal coupling, at zero temperature, at both one and two loop level. This cannot be achieved by the linear curvature term alone and the result remains valid for a very large range of renormalisation scale. Also, we have obtained this result with a cosmological constant as tiny as its current observed value.
@@ 2 The basic set up
$S=-\int d^{d}x\sqrt{-g}\left[\frac{1}{2}g^{\mu\nu}(\nabla_{\mu}\Phi)(\nabla_{\nu}\Phi)+\frac{1}{2}m^{2}\Phi^{2}+\frac{\eta\Phi^{3}}{3!}+\frac{\lambda\Phi^{4}}{4!}+\frac{1}{2}\xi R\Phi^{2}\right]\qquad\qquad(\lambda>0)$ (1)
@@
$\left(\square+m^{2}+\frac{\eta\Phi}{2}+\frac{\lambda\Phi^{2}}{3!}+\xi R\right)\Phi=0$ (2)
We are interested to do quantum field theory in a neighborhood of a given spacetime point. Accordingly, we expand the metric in the Riemann normal coordinate ($y^{\mu}$) around it [79]
$g_{\mu\nu}=\eta_{\mu\nu}-\frac{1}{3}R_{\mu\alpha\nu\beta}y^{\alpha}y^{\beta}-\frac{1}{6}R_{\mu\alpha\nu\beta;\gamma}y^{\alpha}y^{\beta}y^{\gamma}+\left(-\frac{1}{20}R_{\mu\alpha\nu\beta;\gamma\delta}+\frac{2}{45}R_{\alpha\mu\beta\lambda}R^{\lambda}{}_{\gamma\nu\delta}\right)y^{\alpha}y^{\beta}y^{\gamma}y^{\delta}\cdots$
$|g|=1-\frac{1}{3}R_{\alpha\beta}y^{\alpha}y^{\beta}-\frac{1}{6}R_{\alpha\beta;\gamma}y^{\alpha}y^{\beta}y^{\gamma}+\left(\frac{1}{18}R_{\alpha\beta}R_{\gamma\delta}-\frac{1}{90}R^{\lambda}{}_{\alpha\beta}\,^{\kappa}R_{\lambda\gamma\delta\kappa}-\frac{1}{20}R_{\alpha\beta;\gamma\delta}\right)y^{\alpha}y^{\beta}y^{\gamma}y^{\delta}\cdots$ (3)
@@
$G(x,x^{\prime})=g^{-{\frac{1}{4}}}(x)\bar{G}(x,x^{\prime})g^{-{\frac{1}{4}}}(x^{\prime})$ (4)
@@
$\bar{G}(x,x^{\prime})=\int\frac{d^{d}x}{(2\pi)^{d}}e^{ik\cdot y}\bar{G}(k)$ (5)
@@
$\bar{G_{2}}(k)=\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2})^{2}},\qquad\qquad\bar{G_{3}}(k)=\frac{i\left(\frac{1}{6}-\xi\right)(\nabla_{\alpha}R)}{k^{2}+m^{2}}\partial^{\alpha}_{k}\frac{1}{k^{2}+m^{2}}$
$\bar{G_{4}}(k)=\frac{\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}}{(k^{2}+m^{2})^{3}}+\frac{a_{\alpha\beta}}{k^{2}+m^{2}}\partial^{\alpha}_{k}\partial^{\beta}_{k}\frac{1}{k^{2}+m^{2}}$ (8)
@@
$a_{\alpha\beta}=\frac{1}{2}\left(\xi-\frac{1}{6}\right)\nabla_{\alpha}\nabla_{\beta}R+\frac{1}{120}\nabla_{\alpha}\nabla_{\beta}R-\frac{1}{40}\square R_{\alpha\beta}+\frac{1}{30}R_{\alpha}\,^{\lambda}R_{\lambda\beta}-\frac{1}{60}{R^{\rho}\,_{\alpha}}^{\lambda}\,_{\beta}R_{\rho\lambda}-\frac{1}{60}R_{\lambda\mu\rho\alpha}R^{\lambda\mu\rho}{}_{\beta}$ (9)
We shall restrict our calculation to the quadratic order of the curvature only. Up to this order we collect all the terms to write
@@
$+\left[\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}-\frac{2}{3}a^{\lambda}\,_{\lambda}\right]\frac{1}{(k^{2}+m^{2})^{3}}+{\cal O}(R^{3})$ (10)
The above equation will serve as the key ingredient for our computations appearing below.
@@ 3 One loop effective potential at zero temperature
The computation of the divergent part and subsequent renormalisation of the one loop effective action in a curved spacetime can be seen in, e.g. [4] and references therein. In this section we wish to compute below the finite part of the effective potential at zero temperature after renormalisation, up to quadratic order in the curvature. Although the result appearing in this section can easily be found by directly using the standard ‘$\ln{\rm det}$’ formula for the one loop effective action, we wish to briefly sketch below the use of the local momentum space, for the sake of completeness.
@@
$m^{2}_{1}=m^{2}+\frac{1}{2}\lambda\phi^{2}+\eta\phi$
@@
$V_{\rm{eff}}^{1-{\rm loop}}=\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m^{2}_{1})=-\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\ln G(k,m_{1})$ (11)
where $G(k,m_{1})$ is the propagator of the scalar in the flat spacetime with effective mass $m_{1}$. We now replace the flat space propagator by that of the curved spacetime, 10, to find
@@
$=\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\left[\ln(k^{2}+m^{2}_{1})-\frac{(\frac{1}{6}-\xi)R}{(k^{2}+m_{1}^{2})}-\frac{2f_{2}}{(k^{2}+m_{1}^{2})^{2}}+\frac{1}{2}\frac{(\frac{1}{6}-\xi)^{2}R^{2}}{(k^{2}+m_{1}^{2})^{2}}\right]+{\cal O}(R^{3})$ (12)
@@
and $a_{\mu\nu}$ is given by 9. The first term on the right hand side of 12 is the flat space term, whereas the rest are the leading curvature corrections. For $d=4-\epsilon$, the standard momentum integrals appearing above are given by
$\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m_{1}^{2})=\frac{m^{4}_{1}}{(4\pi)^{2}}\left[-\frac{\mu^{-\epsilon}}{\epsilon}+\frac{1}{2}\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\frac{1}{2}-\frac{1}{2}\psi(2)\right]+{\cal O}(\epsilon)$
$-\left(\frac{1}{6}-\xi\right)R\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})}=-\left(\frac{1}{6}-\xi\right)\frac{Rm^{2}_{1}}{(4\pi)^{2}}\left[-\frac{2\mu^{-\epsilon}}{\epsilon}+\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\psi(2)\right]+{\cal O}(\epsilon)$
$\left(\frac{1}{2}{\left(\frac{1}{6}-\xi\right)^{2}R^{2}}-2f_{2}\right)\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})^{2}}=\frac{{\left(\frac{1}{6}-\xi\right)^{2}R^{2}-4f_{2}}}{2(4\pi)^{2}}\left[\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(1)-\ln\frac{m^{2}_{1}}{4\pi\mu^{2}}\right]+{\cal O}(\epsilon)$ (13)
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is the digamma function and $\mu$ is a renormalisation scale. Substituting the three above integrals into 12 yields the (unrenormalised) one loop effective action. The renormalisation of the effective potential or the action can be seen in [4]. The divergent part of the effective potential reads (with $\hbar=1$)
@@
$\left.+\left(\xi-\frac{1}{6}\right)\eta R\phi+\frac{1}{2}\left(m^{2}\lambda+\eta^{2}\right)\phi^{2}+\frac{\lambda^{2}\phi^{4}}{8}+\frac{\eta\lambda\phi^{3}}{2}+\frac{\lambda}{2}\left(\xi-\frac{1}{6}\right)R\phi^{2}-\frac{1}{3}\left(\xi-\frac{1}{5}\right)\square R\right]$ (14)
We may ignore the last term as it is a total divergence. In order to renormalise the divergence associated with the pure curvature terms, we take the gravitational counterterms in the Lagrangian density
@@
$\delta\Lambda=\frac{\mu^{-\epsilon}m^{4}}{32\pi^{2}\epsilon},\qquad\delta\kappa=\frac{\mu^{-\epsilon}m^{2}\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\alpha_{1}=\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\frac{1}{90}$
$\delta\alpha_{2}=-\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\frac{1}{90},\qquad\delta\alpha_{3}=\frac{\mu^{-\epsilon}}{32\pi^{2}\epsilon}\left(\frac{1}{36}+\xi^{2}-\frac{\xi}{3}\right)\qquad$ (16)
The remaining divergence of 14 is absorbed by a matter counterterm Lagrangian
@@
$\delta m^{2}=-\frac{\mu^{-\epsilon}\left(m^{2}\lambda+\eta^{2}\right)}{16\pi^{2}\epsilon},\qquad\delta\xi=-\frac{\mu^{-\epsilon}\lambda\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\tau=-\frac{\mu^{-\epsilon}m^{2}\eta}{16\pi^{2}\epsilon}$
@@
$\delta\gamma=-\frac{\mu^{-\epsilon}\eta\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\eta=-\frac{3\mu^{-\epsilon}\eta\lambda}{16\pi^{2}\epsilon},\qquad\delta\lambda=-\frac{3\mu^{-\epsilon}\lambda^{2}}{16\pi^{2}\epsilon}$ (18)

### v2

@@ Abstract
We revisit the problem of spontaneous symmetry breaking (SSB), its restoration, and phase transition for a self interacting quantum scalar field in a general curved background, at zero and finite temperature. To the best of our knowledge, most of the earlier computations in this context have been done in the linear order in curvature, which may not be very suitable for the Ricci flat spacetimes. One of our objectives is to see whether the higher order terms can bring in qualitatively new physical effects, and thereby attempting to fill in this gap in the literature. We use Bunch and Parker’s local momentum space representation of the Schwinger-DeWitt expansion of the Feynman propagator. Such expansion, being based upon the local Lorentz symmetry of spacetime, essentially probes the leading curvature correction to short scale, ultraviolet quantum processes. We compute the renormalised, background spacetime curvature (up to quadratic order) and temperature dependent one loop effective potential for $\phi^{4}$ plus $\phi^{3}$ self interaction. In particular for the de Sitter spacetime, we have shown for the $\phi^{4}$-theory that we can have SSB even with a positive rest mass squared and positive non-minimal coupling, at zero temperature. This cannot be achieved by the linear curvature term alone and the result remains valid for a very large range of renormalisation scale. Such SSB will generate a field mass that depends upon the spacetime curvature as well as the non-minimal coupling. For a phase transition, we have computed the leading curvature correction to the critical temperature. At finite temperature, symmetry restoration is demonstrated. We also extend some of the above results to two loop level. The symmetry breaking in de Sitter at two loop remains present. We have further motivated the necessity of treating this problem non-perturbatively in some instances.
@@ 1 Introduction
The motivation of studying quantum field theory in curved backgrounds is to investigate the effects of spacetime curvature or geometry on the fluctuations or propagation of quantum fields living in it. This much cultivated yet still active subject, originally supposed to be a prelude to a full theory of quantum gravity, has led to interesting physical predictions like the particle creation in cosmological or black hole backgrounds, fluctuations in the early inflationary universe, regularisation of the stress energy momentum tensor and so on, over decades. We refer our reader to [1, 2, 3, 5, 6] and references therein for a vast historic overview.
@@
We are interested in looking into a self-interacting scalar field theory at zero as well as at finite temperature in a general, regular curved background in this paper. Thermal fields in a curved background may exist in various different scenarios. For instance, the thermal fields living in the early universe, or the hot plasma residing in a star or accreting onto a black hole, the blackbody radiation emitted by a black hole, and so on.
@@
An important aspect of quantum theory of a scalar field is the existence of an effective action and a corresponding effective potential. Such potential arises due to the integration over the quantum fluctuations and it adds up to the tree level self interaction potential. The total effective potential (i.e., the tree plus the loop correction) can be qualitatively different in shape from that of the tree level. Such effective potential may lead to the highly interesting physical phenomenon of spontaneous symmetry breaking, developed in the seminal works in [7, 8, 9, 10]. We further refer our reader to [11, 12, 13, 14, 15, 16, 17, 18] and references therein for various aspects associated with symmetry breaking and effective potentials. Computation of effective actions and potentials for field theories at finite temperature is also a physically well motivated problem e.g. [9, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Interestingly, the effective potential at high temperature ceases to have any symmetry breaking feature even if at zero temperature it has so, known as the symmetry restoration phenomenon. We further refer our reader to [31, 32, 33] and references therein for vast discussion on these issues.
@@
We do not have the scope here to review in detail the developments of quantum field theory in curved spacetimes, and the effective action formalism. We refer our reader to [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] and references therein for some aspects of propagators, renormalisation, effective action and symmetry breaking issues at zero temperature. For thermal propagators and finite temperature effects in flat and curved spacetimes, we refer our reader to [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85] and also references therein. As far as symmetry breaking and phase transition in a general curved background is concerned, most of the earlier works use linear in the curvature approximation and then employ some renormalisation group technique for resummation (e.g. [39, 41, 44, 48] and also references therein). For example, in [44], a renormalisation group derivation of the effective potential for a massless scalar coupled to gauge field, and a discussion on phase transition can be seen. Even though this is a powerful method, it does not work for the Ricci flat spacetimes. In particular, the $R\phi^{2}$-type non-minimal coupling does not contribute to the effective potential for such spacetimes. At finite temperature on the other hand, the Page approximation technique can be useful for static spacetimes [70]. To the best of our knowledge, the thermal effective Lagrangian density in a general curved spacetime for $\phi^{4}$ theory, up to quadratic order in the curvature was first derived in [74], by employing a proper time expansion technique.
@@
The rest of the paper is organised as follows. In the next section, we discuss very briefly the basic technical framework we work in. In 3, we briefly derive the one loop effective potential at zero temperature, up to the quadratic order of the spacetime curvature for quartic and cubic self interactions. In 4, we extend these results to finite temperature, with high temperature approximation. Although as we mentioned, the above problem was first addressed in [74], our notation and the choice of the renormalisation scale will be a bit different. Due to this reason, we shall keep briefly the computations, for the sake of completeness and clarity. We discuss spontaneous symmetry breaking, its restoration and phase transition at zero and finite temperatures in 5, 5.1. We next generalise some of these results to two loop in 6. In particular, the contribution from the two loop double bubble diagram under local approximation to the effective potential, at finite temperature has been explicitly presented, 6.1. Finally, we conclude in 7 with some future motivation for non-perturbative calculations. For the de Sitter spacetime in particular, we have shown that we can have symmetry breaking for a scalar even with positive rest mass squared and a positive non-minimal coupling, at zero temperature, at both one and two loop level. This cannot be achieved by the linear curvature term alone and the result remains valid for a very large range of renormalisation scale.
@@ 2 The basic set up
$S=-\int d^{d}x\sqrt{-g}\left[\frac{1}{2}g^{\mu\nu}(\nabla_{\mu}\Phi)(\nabla_{\nu}\Phi)+\frac{1}{2}m^{2}\Phi^{2}+\frac{\eta\Phi^{3}}{3!}+\frac{\lambda\Phi^{4}}{4!}+\frac{1}{2}\xi R\Phi^{2}\right]\qquad\qquad(\lambda>0).$ (1)
@@
$\left(-\square+m^{2}+\frac{\eta\Phi}{2}+\frac{\lambda\Phi^{2}}{3!}+\xi R\right)\Phi=0.$ (2)
We are interested to do QFT in a neighborhood of a given spacetime point. Accordingly, we expand the metric in the Riemann normal coordinate ($y^{\mu}$) around it (see e.g., [86] for technical detail and also for some original references)
$g_{\mu\nu}=\eta_{\mu\nu}-\frac{1}{3}R_{\mu\alpha\nu\beta}y^{\alpha}y^{\beta}-\frac{1}{6}R_{\mu\alpha\nu\beta;\gamma}y^{\alpha}y^{\beta}y^{\gamma}+\left(-\frac{1}{20}R_{\mu\alpha\nu\beta;\gamma\delta}+\frac{2}{45}R_{\alpha\mu\beta\lambda}R^{\lambda}{}_{\gamma\nu\delta}\right)y^{\alpha}y^{\beta}y^{\gamma}y^{\delta}\cdots,$
$|g|=1-\frac{1}{3}R_{\alpha\beta}y^{\alpha}y^{\beta}-\frac{1}{6}R_{\alpha\beta;\gamma}y^{\alpha}y^{\beta}y^{\gamma}+\left(\frac{1}{18}R_{\alpha\beta}R_{\gamma\delta}-\frac{1}{90}R^{\lambda}{}_{\alpha\beta}\,^{\kappa}R_{\lambda\gamma\delta\kappa}-\frac{1}{20}R_{\alpha\beta;\gamma\delta}\right)y^{\alpha}y^{\beta}y^{\gamma}y^{\delta}\cdots.$ (3)
@@
$G(x,x^{\prime})=g^{-{\frac{1}{4}}}(x)\bar{G}(x,x^{\prime})g^{-{\frac{1}{4}}}(x^{\prime}).$ (4)
@@
$\bar{G}(x,x^{\prime})=\int\frac{d^{d}x}{(2\pi)^{d}}e^{ik\cdot y}\bar{G}(k).$ (5)
@@
$\bar{G_{2}}(k)=\frac{\left(\frac{1}{6}-\xi\right)R}{(k^{2}+m^{2})^{2}},\qquad\qquad\bar{G_{3}}(k)=\frac{i\left(\frac{1}{6}-\xi\right)(\nabla_{\alpha}R)}{k^{2}+m^{2}}\partial^{\alpha}_{k}\frac{1}{k^{2}+m^{2}},$
$\bar{G_{4}}(k)=\frac{\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}}{(k^{2}+m^{2})^{3}}+\frac{a_{\alpha\beta}}{k^{2}+m^{2}}\partial^{\alpha}_{k}\partial^{\beta}_{k}\frac{1}{k^{2}+m^{2}},$ (8)
@@
$a_{\alpha\beta}=\frac{1}{2}\left(\xi-\frac{1}{6}\right)\nabla_{\alpha}\nabla_{\beta}R+\frac{1}{120}\nabla_{\alpha}\nabla_{\beta}R-\frac{1}{40}\square R_{\alpha\beta}+\frac{1}{30}R_{\alpha}\,^{\lambda}R_{\lambda\beta}-\frac{1}{60}{R^{\rho}\,_{\alpha}}^{\lambda}\,_{\beta}R_{\rho\lambda}-\frac{1}{60}R_{\lambda\mu\rho\alpha}R^{\lambda\mu\rho}{}_{\beta}.$ (9)
We shall restrict our calculation up to the quadratic order of the curvature. We next collect all such terms to write a local momentum space representation of the Schwinger-DeWitt series expansion of the Feynman propagator,
@@
$+\left[\left(\frac{1}{6}-\xi\right)^{2}{R^{2}}-\frac{2}{3}a^{\lambda}\,_{\lambda}\right]\frac{1}{(k^{2}+m^{2})^{3}}+{\cal O}(R^{3}).$ (10)
Note that the above propagator is locally Lorentz invariant, by the virtue of 3. The above equation will serve as the key ingredient for our computations appearing below.
@@ 3 One loop effective potential at zero temperature
The computation of the divergent part and subsequent renormalisation of the one loop effective action in a curved spacetime can be seen in, e.g. [6] and references therein. In this section, we wish to compute below the finite part of the effective potential at zero temperature after renormalisation, up to quadratic order in the curvature. Although the result appearing in this section can easily be found by directly using the standard ‘$\ln{\rm det}$’ formula for the one loop effective action, we wish to briefly sketch below the use of the local momentum space, for the sake of completeness.
@@
$m^{2}_{1}=m^{2}+\frac{1}{2}\lambda\phi^{2}+\eta\phi.$ (11)
Clearly, for the discussion of SSB, the cubic self interaction is not relevant. However, we wish to keep the same in our general derivation for the effective potential, firstly because the $\lambda\phi^{4}+\eta\phi^{3}$ theory is a general renormalisable model in $(3+1)$-dimensions and second, the antisymmetric cubic self-interaction complements the symmetric and bounded nature of the quartic one. Most importantly, there can be various instances other than SSB where the asymmetry of the $\lambda\phi^{4}+\eta\phi^{3}$ self interaction could bring in some interesting physical effects.
@@
$V_{\rm{eff}}^{1-{\rm loop}}=-\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\ln G_{\rm{Min}}(k,m_{1}),$ (12)
where $G_{\rm{Min}}=(k^{2}+m_{1}^{2})^{-1}$ is the propagator of the scalar in the flat spacetime with an effective mass squared $m_{1}^{2}$, 11. We now replace the flat space propagator by that of the curved spacetime, 10, to find
@@
$=\frac{\hbar}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\left[\ln(k^{2}+m^{2}_{1})-\frac{(\frac{1}{6}-\xi)R}{(k^{2}+m_{1}^{2})}-\frac{2f_{2}}{(k^{2}+m_{1}^{2})^{2}}+\frac{1}{2}\frac{(\frac{1}{6}-\xi)^{2}R^{2}}{(k^{2}+m_{1}^{2})^{2}}\right]+{\cal O}(R^{3}),$ (13)
@@
and $a_{\mu\nu}$ is given by 9. The first term on the right hand side of 13 is the flat space term, whereas the rest are the leading curvature corrections. For $d=4-\epsilon$, the three standard momentum integrals appearing above are given by
$\int\frac{d^{d}k}{(2\pi)^{d}}\ln(k^{2}+m_{1}^{2})=\frac{m^{4}_{1}}{(4\pi)^{2}}\left[-\frac{\mu^{-\epsilon}}{\epsilon}+\frac{1}{2}\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\frac{1}{4}-\frac{1}{2}\psi(2)\right]+{\cal O}(\epsilon),$
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})}=\frac{m^{2}_{1}}{(4\pi)^{2}}\left[-\frac{2\mu^{-\epsilon}}{\epsilon}+\ln\frac{m_{1}^{2}}{4\pi\mu^{2}}-\psi(2)\right]+{\cal O}(\epsilon),$
$\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2}+m_{1}^{2})^{2}}=\frac{1}{(4\pi)^{2}}\left[\frac{2\mu^{-\epsilon}}{\epsilon}+\psi(1)-\ln\frac{m^{2}_{1}}{4\pi\mu^{2}}\right]+{\cal O}(\epsilon),$ (14)
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is the digamma function and $\mu$ is a renormalisation scale. Substituting the three above integrals into 13 yields the (unrenormalised) one loop effective action. The renormalisation of the effective potential or the action can be seen in [6]. The purely divergent part of the effective potential reads (with $\hbar=1$)
@@
$\left.+\left(\xi-\frac{1}{6}\right)\eta R\phi+\frac{1}{2}\left(m^{2}\lambda+\eta^{2}\right)\phi^{2}+\frac{\lambda^{2}\phi^{4}}{8}+\frac{\eta\lambda\phi^{3}}{2}+\frac{\lambda}{2}\left(\xi-\frac{1}{6}\right)R\phi^{2}-\frac{1}{3}\left(\xi-\frac{1}{5}\right)\square R\right].$ (15)
We may ignore the last term as it is a total divergence. However note that, since 13 and 15 contain $\phi$-independent terms, probably one should not regard them as effective potential. In a strict sense, they should instead be attributed as effective Lagrangian density minus the kinetic term.
In order to renormalise the divergence associated with the pure curvature terms of 15, we take the gravitational counterterms in the Lagrangian density
@@
$\delta\Lambda=\frac{\mu^{-\epsilon}m^{4}}{32\pi^{2}\epsilon},\qquad\delta\kappa=\frac{\mu^{-\epsilon}m^{2}\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\alpha_{1}=\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\frac{1}{90},$
$\delta\alpha_{2}=-\frac{\mu^{-\epsilon}}{16\pi^{2}\epsilon}\frac{1}{90},\qquad\delta\alpha_{3}=\frac{\mu^{-\epsilon}}{32\pi^{2}\epsilon}\left(\frac{1}{36}+\xi^{2}-\frac{\xi}{3}\right).\qquad$ (17)
The remaining divergence of 15 is absorbed by a matter counterterm Lagrangian
@@
$\delta m^{2}=-\frac{\mu^{-\epsilon}\left(m^{2}\lambda+\eta^{2}\right)}{16\pi^{2}\epsilon},\qquad\delta\xi=-\frac{\mu^{-\epsilon}\lambda\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\tau=-\frac{\mu^{-\epsilon}m^{2}\eta}{16\pi^{2}\epsilon},$
@@
$\delta\gamma=-\frac{\mu^{-\epsilon}\eta\left(\xi-\frac{1}{6}\right)}{16\pi^{2}\epsilon},\qquad\delta\eta=-\frac{3\mu^{-\epsilon}\eta\lambda}{16\pi^{2}\epsilon},\qquad\delta\lambda=-\frac{3\mu^{-\epsilon}\lambda^{2}}{16\pi^{2}\epsilon}.$ (19)